# Inertial-Range Suppression and Ponderomotive Density Cavitation in Broadband Sub-Alfvénic Turbulence under Plasma Sheet Boundary Layer Conditions

### 15 Mar 2026 v1

@@ Abstract
Kinetic Alfvén waves (KAWs) are among the most pervasive electromagnetic fluctuations in magnetized astrophysical plasmas, from Earth’s magnetospheric boundary layers to the turbulent intracluster medium of galaxy clusters. Their ponderomotive coupling to compressive density fluctuations remains incompletely understood in the broadband turbulent regime. We present two-dimensional pseudospectral simulations of the modified nonlinear Schrödinger–magnetosonic (MNLS–MS) system governing KAW envelopes, initialized with a broadband power-law spectrum ($|\psi(\mathbf{k})|^{2}\propto k^{-5/6}$) spanning many interacting modes, at $\beta\sim 0.1$$0.3$ representative of plasma sheet boundary layer (PSBL) conditions. A fourth-order Runge–Kutta scheme on a $256\times 256$ grid integrates the system to $t=40$ (normalized), with total energy conserved to within $0.085\%$. The nonlinearity parameter $\chi_{\mathrm{NL}}\approx 0.25$ confirms broadband sub-Alfvénic turbulence throughout. Magnetic field intensity and plasma density develop spatially intermittent, filamentary structures within the first few wave periods, consistent with ponderomotive density cavitation and plasma expulsion from wave-intense regions. The magnetic energy spectra show inertial-range suppression, with a rapid transition from injection ($k<0.3$) to dissipation without an extended power-law cascade, in agreement with the moderate magnetic Reynolds number ($\mathrm{Re}\sim 250$$370$) of the simulation and the observationally constrained range for PSBL turbulence. These results provide numerical evidence that broadband KAW turbulence self-organizes into coherent density structures at kinetic scales, and that the spectral character of such turbulence is governed primarily by moderate-Reynolds-number constraints rather than by the wave physics alone.
Corresponding authors:
Mani K Chettri: mkchettri8@gmail.com
Hemam D. Singh: hemam.singh@nsut.ac.in
Keywords: kinetic Alfvén waves; modified nonlinear Schrödinger equation; plasma turbulence; ponderomotive force; density cavities; pseudospectral method; plasma sheet boundary layer; astrophysical plasmas
@@ 1 Introduction
Alfvénic turbulence is a universal feature of magnetized astrophysical plasmas, mediating the transfer of energy from the large fluid scales at which it is injected down to the kinetic scales of individual particles, where it is ultimately thermalized. At perpendicular spatial scales approaching the ion gyroradius $\rho_{i}$ or the ion inertial length $d_{i}$, ideal shear Alfvén waves acquire dispersive corrections and transform into kinetic Alfvén waves (KAWs), a transition that is observationally well established in the solar wind (Leamon et al., 1999; Bale et al., 2005; Chen et al., 2013; Sahraoui et al., 2009), in Earth’s magnetosphere (Chaston et al., 2008; Wygant and others, 2002), and in the magnetosheath (Roberts and others, 2018; Gershman and others, 2017b). KAWs carry a finite parallel electric field $E_{\parallel}$ that opens a direct channel for wave-particle energy exchange through Landau damping (Howes et al., 2008; Chen et al., 2019), and at finite amplitude they exert a ponderomotive force on the bulk plasma, a low-frequency, gradient-driven pressure $\propto{-}\nabla|\delta B_{y}|^{2}/8\pi$, which expels ions from regions of high wave intensity and forms localized density depletions. These two effects, particle heating and density structuring, make KAWs important agents not only in Earth’s magnetotail (Wygant and others, 2002; Stawarz and others, 2017; Chettri et al., 2026) but also in the broader context of astrophysical plasma heating. In solar coronal loops, parametric decay of standing Alfvén waves generates compressive daughter modes with observationally testable signatures (Terradas et al., 2022). In the inner heliosphere, coupled KAW and ion-acoustic wave dynamics have been studied using Parker Solar Probe data (Chettri et al., 2024). In the interstellar medium, kinetic-scale Alfvénic turbulence shapes the scattering properties of pulsars and governs cosmic-ray transport (Lazarian and Beresnyak, 2006; Yan and Lazarian, 2004). In the turbulent intracluster medium of galaxy clusters, where magnetic fields thread hot, weakly collisional gas, analogous wave-plasma interactions on scales below the Coulomb mean free path have been invoked to explain observed magnetic field tangling and sub-Larmor density fluctuations (Brunetti and Lazarian, 2007; Schekochihin and others, 2009).
@@
The theoretical framework for KAW nonlinear dynamics rests on foundational studies spanning several decades. Hasegawa and Chen (1976) identified parametric decay as the primary instability by which a monochromatic KAW transfers energy to compressive daughter modes. Shukla and Stenflo (1999) derived the coupled KAW–magnetosonic equations that describe ponderomotive density cavity formation in the quasi-static limit and demonstrated the linear anti-correlation $n/n_{0}\propto-|\delta B_{y}|^{2}$. More recently, three-dimensional hybrid simulations of circularly polarized Alfvén waves by Verscharen and others (2024) have confirmed that the parametric decay instability operates across a broad spectrum of wave vectors. Goldreich and Sridhar (1995) and Boldyrev and Perez (2012) established the theoretical spectral scalings expected when Alfvénic turbulence reaches a critically balanced cascade, providing benchmarks against which kinetic-scale simulations can be compared. Our group recently derived and validated a modified nonlinear Schrödinger equation (MNLSE) framework that self-consistently incorporates KAW dispersion, ponderomotive nonlinearity, and phenomenological Landau damping, and tested it against MMS magnetosheath observations (Chettri et al., 2025). Separately, Chettri et al. (2026) reported direct MMS observations of KAW turbulence signatures in the outer PSBL, establishing the observational context for the present numerical study. Nevertheless, the ponderomotive scaling relation has been tested chiefly in analytical or single-wave-packet settings; how it behaves within a broadband fluctuation environment where multiple interacting modes compete and interfere remains much less well characterized.
@@
This gap motivates the present work. We solve the coupled MNLS–magnetosonic system of Chettri et al. (2025) in two spatial dimensions using a pseudospectral method on a $256\times 256$ grid, initialized with a broadband turbulent spectrum representative of PSBL conditions. The PSBL provides a well-observed astrophysical boundary layer where magnetic energy, particle fluxes, and broadband electromagnetic turbulence coexist at directly measurable intensities, giving concrete observational anchors for the simulation parameters. Our goals are fourfold: (1) to examine how KAW ponderomotive coupling operates when the driving is turbulent rather than monochromatic; (2) to characterize the spatial structure and temporal persistence of the resulting density fluctuations; (3) to assess the spectral character of the magnetic energy cascade at moderate Reynolds numbers; and (4) to provide diagnostics (energy conservation, nonlinearity parameter, energy partition) that allow meaningful comparison with in-situ observations and theoretical expectations.
@@ 2.1 Modified Nonlinear Schrödinger Equation
Following Chettri et al. (2025), the slowly-varying envelope $\psi$ satisfies the modified nonlinear Schrödinger (MNLS) equation
@@
Because $Q>0$, regions of enhanced $|\psi|^{2}$ expel plasma and decrease the local refractive index, reinforcing the growth of filamentary intensity peaks through ponderomotive self-focusing (Shukla and Stenflo, 1999; Hasegawa and Chen, 1976).
The Landau damping coefficient $\Gamma$ approximates the collisionless electron (and ion) damping of the KAW envelope. Following Chettri et al. (2025), it is evaluated from the imaginary part of the linearized kinetic dispersion function at $(k_{0},\omega_{0})$:
$\Gamma=\frac{|\,\mathrm{Im}[\omega_{\mathrm{kin}}(k_{0})]|}{\omega_{0}}\,\omega_{0}.$ (4)
In the low-$\beta$ PSBL regime ($\beta\sim 0.1$$0.3$, $T_{i}/T_{e}=2$), resonant damping is well below the wave frequency, so $\Gamma$ acts as a perturbative correction to the dispersive and nonlinear dynamics. It is retained for consistency with the high-$\beta$ magnetosheath analysis of Chettri et al. (2025); the full kinetic treatment of velocity-space resonances is deferred to gyrokinetic approaches such as those described in Schekochihin and others (2009) and Howes et al. (2008).
@@ 2.2 Ponderomotively Driven Density Equation
The ponderomotive force $\mathbf{f}_{\mathrm{pond}}\propto-\nabla|\psi|^{2}$ drives compressive fluctuations through the magnetosonic branch. Following Shukla and Stenflo (1999) and Chettri et al. (2025), the general form of the density equation is
@@
$\left(\frac{\partial^{2}}{\partial t^{2}}-C_{s}^{2}\,\frac{\partial^{2}}{\partial z^{2}}\right)\frac{n}{n_{0}}=\Gamma_{2}\,\frac{\partial^{2}}{\partial z^{2}}|\psi|^{2},$ (6)
with $\Gamma_{2}$ as defined in equation (3). This approximation is appropriate for the low-$\beta$ conditions explored here, where the acoustic speed satisfies $C_{s}\ll v_{A}$ and the perpendicular pressure balance is maintained by the ambient magnetic field rather than by the thermal plasma pressure. Equations (1) and (6) form the coupled MNLS–MS system solved in the present simulations.
@@
The density perturbations generated by equation (6) represent quasi-modes continuously maintained by the KAW ponderomotive source, rather than freely propagating ion acoustic waves. In the magnetotail, where $T_{i}/T_{e}\sim 1$33, ion Landau damping strongly suppresses free acoustic propagation (Verscharen and others, 2024; Terradas et al., 2022). Energy transferred from the KAW envelope into these density quasi-modes can thermalize rapidly via ion Landau damping, providing a direct ion heating channel not resolved by the present fluid model but important for the astrophysical interpretation.
@@ 3.1 Algorithm
The MNLS–MS system, equations (1) and (6), is solved in two spatial dimensions ($x$, $z$) using a pseudospectral method following the approach of Chettri et al. (2025). All spatial derivatives are computed in Fourier space via the Fast Fourier Transform (FFT), which provides spectral accuracy for smooth fields. Time integration uses a classical fourth-order Runge–Kutta (RK4) scheme with time step $\Delta t=10^{-5}$ (normalized).
@@ 3.2 Simulation Parameters
Table 1 lists all simulation parameters. Plasma parameters are chosen to represent the near-Earth PSBL at $X_{\mathrm{GSM}}\approx-7\,R_{E}$, based on observational compilations from Stawarz and others (2017) and statistical analyses in Du and others (2011) and Zhou and others (2012). The values $B_{0}=63$ nT, $n_{0}=0.72$ cm-3, and $T_{e}=2.6$ keV are taken directly from observations, and the temperature ratio $T_{i}/T_{e}=2$ is consistent with Hydra spacecraft measurements in this region (Shrivastava et al., 2015). Together these yield $\beta\sim 0.1$$0.3$, placing the simulation in the warm plasma regime where electromagnetic KAW generation and propagation are most efficient (Denton and others, 2010).
@@
Table 1: Simulation parameters. Observed plasma values $B_{0}$, $n_{0}$, and $T_{e}$ are from Stawarz and others (2017); $T_{i}/T_{e}$ from Shrivastava et al. (2015).
@@ 4.1 Energy Conservation and Nonlinearity
Figure 1 shows three global diagnostics tracked throughout the run. Panel (a) displays the total energy $E_{\mathrm{tot}}$, the sum of magnetic (KAW), kinetic (wave), and acoustic (density) components integrated over the domain. Over the full integration to $t=40$, the energy drifts by only $0.085\%$, monotonically and without runaway behaviour, confirming that the hyperviscosity removes energy at grid scales rather than allowing it to accumulate at physical ones.
Panel (b) shows the partition between magnetic and density energy on a logarithmic scale. The magnetic component dominates throughout at a roughly constant normalized value near $0.25$; the acoustic component sits near $10^{-3}$ and declines slowly. This two-order-of-magnitude gap is physically meaningful: because energy transferred from KAWs to compressive modes scales as $(n/n_{0})^{2}$, and the density response remains small at the explored amplitudes, the acoustic component necessarily carries a minor share of the total budget. This does not imply weak ponderomotive coupling; the density structures in Section 4.3 are spatially prominent. It reflects the dominance of the Alfvénic energy reservoir.
Panel (c) shows the nonlinearity parameter $\chi_{\mathrm{NL}}=\sigma(|\delta B_{y}|\,|n/n_{0}|)\,/\,\sigma(|\delta B_{y}|)$, which measures the relative strength of the ponderomotive coupling term against the linear envelope amplitude. The values remain in the range 0.20.20.280.28 throughout, well below the critical threshold $\chi_{\mathrm{NL}}=1$ at which the transition to strong turbulence occurs. The simulation therefore operates in the sub-Alfvénic, moderately nonlinear regime characteristic of quieter PSBL intervals and of many other astrophysical boundary layers (Galtier et al., 2000; Matthaeus and Velli, 2011).
Figure 1: Global diagnostic quantities over the full simulation. (a) Total energy evolution showing a drift of only $0.085\%$, confirming numerical reliability. (b) Energy partition between magnetic (blue) and density (green) components on a logarithmic scale; the magnetic component dominates by roughly two orders of magnitude throughout. (c) Nonlinearity parameter $\chi_{\mathrm{NL}}$ as a function of time; values near $0.25$, well below the strong-turbulence threshold $\chi_{\mathrm{NL}}=1$ (red dashed line), confirm sub-Alfvénic operation.

### 18 Apr 2026 v2

@@ Abstract
Kinetic Alfvén waves (KAWs) are among the most pervasive electromagnetic fluctuations in magnetized astrophysical plasmas, from Earth’s magnetospheric boundary layers to the turbulent intracluster medium of galaxy clusters. Their ponderomotive coupling to compressive density fluctuations remains incompletely understood in the broadband turbulent regime. We present two-dimensional pseudospectral simulations of the modified nonlinear Schrödinger–magnetosonic (MNLS–MS) system governing KAW envelopes, initialized with a broadband power-law spectrum ($|\psi(\mathbf{k})|^{2}\propto k^{-5/6}$) spanning many interacting modes, at $\beta\sim 0.1$$0.3$ representative of plasma sheet boundary layer (PSBL) conditions. A fourth-order Runge–Kutta scheme on a $256\times 256$ grid integrates the system to $t=40$ (normalized), with total energy conserved to within $0.085\%$ in the undamped run; a damped run with dissipation loses $\sim 4\%$ of the magnetic energy over the same interval. The nonlinearity parameter $\chi_{\mathrm{NL}}\approx 0.25$ confirms broadband sub-Alfvénic turbulence throughout. Magnetic field intensity and plasma density develop spatially intermittent, filamentary structures within the first few wave periods, consistent with ponderomotive density cavitation and plasma expulsion from wave-intense regions. The magnetic energy spectra show inertial-range suppression, with a rapid transition from injection ($k<0.3$) to dissipation without an extended power-law cascade, in agreement with the moderate magnetic Reynolds number ($\mathrm{R_{m}}\sim 250$$370$) of the simulation and the observationally constrained range for PSBL turbulence. These results provide numerical evidence that broadband KAW turbulence self-organizes into coherent density structures at kinetic scales, and that the spectral character of such turbulence is governed primarily by moderate-Reynolds-number constraints rather than by the wave physics alone.
keywords:
[…]
@@ 1 Introduction
Alfvénic turbulence is a universal feature of magnetized astrophysical plasmas, mediating the transfer of energy from the large fluid scales at which it is injected down to the kinetic scales of individual particles, where it is ultimately thermalized. At perpendicular spatial scales approaching the ion gyroradius $\rho_{i}$ or the ion inertial length $d_{i}$, ideal shear Alfvén waves acquire dispersive corrections and transform into kinetic Alfvén waves (KAWs), a transition that is observationally well established in the solar wind (leamon1999; bale2005; chen2013; sahraoui2009), in Earth’s magnetosphere (chaston2008; wygant2002), and in the magnetosheath (roberts2018; gershman2017). KAWs carry a finite parallel electric field $E_{\parallel}$ that opens a direct channel for wave-particle energy exchange through Landau damping (howes2008; chen2019), and at finite amplitude they exert a ponderomotive force on the bulk plasma, a low-frequency, gradient-driven pressure $\propto{-}\nabla|\delta B_{y}|^{2}/8\pi$, which expels ions from regions of high wave intensity and forms localized density depletions. These two effects, particle heating and density structuring, make KAWs important agents not only in Earth’s magnetotail (wygant2002; stawarz2017; chettri2025mms) but also in the broader context of astrophysical plasma heating. In solar coronal loops, parametric decay of standing Alfvén waves generates compressive daughter modes with observationally testable signatures (terradas2022). In the inner heliosphere, coupled KAW and ion-acoustic wave dynamics have been studied using Parker Solar Probe data (chettri2024raa). In the interstellar medium, kinetic-scale Alfvénic turbulence shapes the scattering properties of pulsars and governs cosmic-ray transport (lazarian2006; yan2004). In the turbulent intracluster medium of galaxy clusters, where magnetic fields thread hot, weakly collisional gas, analogous wave-plasma interactions on scales below the Coulomb mean free path have been invoked to explain observed magnetic field tangling and sub-Larmor density fluctuations (brunetti2007; schekochihin2009).
@@
The theoretical framework for KAW nonlinear dynamics rests on foundational studies spanning several decades. hasegawa1976parametric identified parametric decay as the primary instability by which a monochromatic KAW transfers energy to compressive daughter modes. shukla1999 derived the coupled KAW–magnetosonic equations that describe ponderomotive density cavity formation in the quasi-static limit and demonstrated the linear anti-correlation $n/n_{0}\propto-|\delta B_{y}|^{2}$. More recently, three-dimensional hybrid simulations of circularly polarized Alfvén waves by verscharen2024 have confirmed that the parametric decay instability operates across a broad spectrum of wave vectors. goldreich1995 and boldyrev2012 established the theoretical spectral scalings expected when Alfvénic turbulence reaches a critically balanced cascade, providing benchmarks against which kinetic-scale simulations can be compared. Our group recently derived and validated a modified nonlinear Schrödinger equation (MNLSE) framework that self-consistently incorporates KAW dispersion, ponderomotive nonlinearity, and phenomenological Landau damping, and tested it against MMS magnetosheath observations (chettri2025damped). Separately, chettri2025mms reported direct MMS observations of KAW turbulence signatures in the outer PSBL, establishing the observational context for the present numerical study. Nevertheless, the ponderomotive scaling relation has been tested chiefly in analytical or single-wave-packet settings; how it behaves within a broadband fluctuation environment where multiple interacting modes compete and interfere remains much less well characterized.
@@
This gap motivates the present work. We solve the coupled MNLS–magnetosonic system of chettri2025damped in two spatial dimensions using a pseudospectral method on a $256\times 256$ grid, initialized with a broadband turbulent spectrum representative of PSBL conditions. The PSBL provides a well-observed astrophysical boundary layer where magnetic energy, particle fluxes, and broadband electromagnetic turbulence coexist at directly measurable intensities, giving concrete observational anchors for the simulation parameters. Our goals are fourfold: (1) to examine how KAW ponderomotive coupling operates when the driving is turbulent rather than monochromatic; (2) to characterize the spatial structure and temporal persistence of the resulting density fluctuations; (3) to assess the spectral character of the magnetic energy cascade at moderate Reynolds numbers; and (4) to provide diagnostics (energy conservation, nonlinearity parameter, and undamped-vs-damped energy comparison) that allow meaningful comparison with in-situ observations and theoretical expectations.
@@ 2.1 Modified Nonlinear Schrödinger Equation
Following chettri2025damped, the slowly-varying envelope $\psi$ satisfies the modified nonlinear Schrödinger (MNLS) equation
@@
Because $Q>0$, regions of enhanced $|\psi|^{2}$ expel plasma and decrease the local refractive index, reinforcing the growth of filamentary intensity peaks through ponderomotive self-focusing (shukla1999; hasegawa1976parametric).
The Landau damping coefficient $\Gamma$ approximates the collisionless electron (and ion) damping of the KAW envelope. Following chettri2025damped, it is evaluated from the imaginary part of the linearized kinetic dispersion function at $(k_{0},\omega_{0})$:
$\Gamma=|\,\mathrm{Im}[\omega_{\mathrm{kin}}(k_{0})]|$ (4)
In the low-$\beta$ PSBL regime ($\beta\sim 0.1$$0.3$, $T_{i}/T_{e}=2$), resonant damping is well below the wave frequency, so $\Gamma$ acts as a perturbative correction to the dispersive and nonlinear dynamics. It is retained for consistency with the high-$\beta$ magnetosheath analysis of chettri2025damped; the full kinetic treatment of velocity-space resonances is deferred to gyrokinetic approaches such as those described in schekochihin2009 and howes2008.
@@ 2.2 Ponderomotively Driven Density Equation
The ponderomotive force $\mathbf{f}_{\mathrm{pond}}\propto-\nabla|\psi|^{2}$ drives compressive fluctuations through the magnetosonic branch. Following shukla1999 and chettri2025damped, the general form of the density equation is
@@
$\left(\frac{\partial^{2}}{\partial t^{2}}-C_{s}^{2}\,\frac{\partial^{2}}{\partial z^{2}}\right)\frac{n}{n_{0}}=\Gamma_{2}\,\frac{\partial^{2}}{\partial z^{2}}|\psi|^{2}.$ (6)
The strong ambient field $\mathbf{B}_{0}$ resists perpendicular plasma displacement, so ponderomotively driven density perturbations propagate preferentially along the field. At the plasma parameters of Table 1, $C_{s}/v_{A}\approx 0.31$, so the omitted perpendicular term $C_{s}^{2}\,\partial^{2}/\partial x^{2}$ is suppressed relative to the retained parallel term by a factor $C_{s}^{2}/v_{A}^{2}\approx 0.09$; the approximation introduces corrections of order $9\%$, within the first-order accuracy of the MNLS–MS framework and consistent with the canonical treatment of shukla1999. The simulations also operate in the quasi-static limit ($\chi_{\mathrm{NL}}\approx 0.25\ll 1$); in this limit the common spatial operator cancels from both sides of equation (5), and the ponderomotive scaling relation (7) is independent of whether $\partial^{2}/\partial z^{2}$ or $\nabla^{2}$ is used. The central anti-correlation results of Section 4.3 are therefore unaffected by this approximation. Inclusion of the full Laplacian is identified as a direction for future work in Section 5.4. Equations (1) and (6) form the coupled MNLS–MS system solved in the present simulations.
@@
The density perturbations generated by equation (6) represent quasi-modes continuously maintained by the KAW ponderomotive source, rather than freely propagating ion acoustic waves. In the magnetotail, where $T_{i}/T_{e}\sim 1$33, ion Landau damping strongly suppresses free acoustic propagation (verscharen2024; terradas2022). Energy transferred from the KAW envelope into these density quasi-modes can thermalize rapidly via ion Landau damping, providing a direct ion heating channel not resolved by the present fluid model but important for the astrophysical interpretation.
@@ 3.1 Algorithm
The MNLS–MS system, equations (1) and (6), is solved in two spatial dimensions ($x$, $z$) using a pseudospectral method following the approach of chettri2025damped. All spatial derivatives are computed in Fourier space via the Fast Fourier Transform (FFT), which provides spectral accuracy for smooth fields. Time integration uses a classical fourth-order Runge–Kutta (RK4) scheme with time step $\Delta t=10^{-5}$ (normalized).
@@ 3.2 Simulation Parameters
Table 1 lists all simulation parameters. Plasma parameters are chosen to represent the near-Earth PSBL at $X_{\mathrm{GSM}}\approx-7\,R_{E}$, based on observational compilations from stawarz2017 and statistical analyses in du2011 and zhou2012. The values $B_{0}=63$ nT, $n_{0}=0.72$ cm-3, and $T_{e}=2.6$ keV are taken directly from observations, and the temperature ratio $T_{i}/T_{e}=2$ is consistent with Hydra spacecraft measurements in this region (shrivastava2015). Together these yield $\beta\sim 0.1$$0.3$, placing the simulation in the warm plasma regime where electromagnetic KAW generation and propagation are most efficient (denton2010).
@@
Table 1: Simulation parameters. Observed plasma values $B_{0}$, $n_{0}$, and $T_{e}$ are from stawarz2017; $T_{i}/T_{e}$ from shrivastava2015.
@@ 4.1 Energy Conservation and Nonlinearity
Figure 1 shows two diagnostic quantities tracked throughout the run. Panel (a) displays the normalized magnetic energy $E_{\mathrm{mag}}(t)/E_{\mathrm{mag}}(0)$ for two cases: an undamped run ($\Gamma=0$, blue solid) and a damped run with the Landau coefficient evaluated from equation (4) (red dashed). In the undamped case the energy is essentially flat, drifting by only $0.085\%$ over the full integration to $t=40$; this drift is attributable entirely to the hyperviscosity $\nu_{h}k^{6}$, acting exclusively above $3k_{\mathrm{max}}/4$, well above the injection scales where most energy resides ($k\lesssim 0.3$). In the damped case, Landau dissipation removes $\sim 4\%$ of the magnetic energy by $t=40$, confirming that the Landau channel produces a measurable but modest energy drain at these parameters. Because the scalar $\Gamma$ is evaluated at a single carrier wavenumber and cannot consistently represent damping across the full broadband spectrum (k=0.08k=0.080.50.5), where the kinetic rate $\gamma(k)$ varies by roughly an order of magnitude, we retain $\Gamma=0$ to verify energy conservation, providing a validation benchmark for the numerical code.
Panel (b) shows the nonlinearity parameter $\chi_{\mathrm{NL}}$ for both the undamped and damped runs, calculated as the ratio of the standard deviation of the product $|\delta B_{y}|\,|n/n_{0}|$ to that of $|\delta B_{y}|$ alone, evaluated over the simulation domain. The two curves are virtually identical, confirming that the sub-Alfvénic regime is robust to the inclusion of Landau damping. The values remain in the range 0.20.20.280.28 throughout, well below the critical threshold $\chi_{\mathrm{NL}}=1$ at which the transition to strong turbulence occurs. The simulation therefore operates in the sub-Alfvénic, moderately nonlinear regime characteristic of quieter PSBL intervals and of many other astrophysical boundary layers (galtier2000; matthaeus2011).
Figure 1: (a) Normalized magnetic energy $E_{\mathrm{mag}}(t)/E_{\mathrm{mag}}(0)$: the undamped run (blue solid, $\Gamma=0$) drifts only $0.085\%$, while the damped run (red dashed) loses $\sim 4\%$ to Landau dissipation, confirming that hyperviscosity and Landau damping act at their expected levels. (b) Nonlinearity parameter $\chi_{\mathrm{NL}}$ as a function of time for both runs; values near $0.25$, well below the strong-turbulence threshold $\chi_{\mathrm{NL}}=1$ (grey dotted line), confirm sub-Alfvénic operation regardless of damping.